# Adaptive Real-Time Multi-Loss Function Optimization Using Dynamic Memory Fusion Framework: A Case Study on Breast Cancer Segmentation

### v1

@@ Abstract
[inst1]organization=Faculty of Electrical & Robotics, Shahrood University of Technology, city=Shahrood, country=Iran
@@ 1 Introduction
In deep learning, the strategic selection, combination, and dynamic weighting of loss functions are crucial for optimizing performance across various tasks. This is particularly important when dealing with challenges like class imbalance, boundary precision, and feature learning. Deep learning models can achieve superior results in various domains by carefully considering these factors and employing appropriate loss functions. In GBE-Net [1], the combination of cross-entropy and dice loss is shown to enhance segmentation accuracy, particularly in preserving the boundaries of lesions. Similarly, MFMSNet integrates dice and binary cross-entropy losses to handle class imbalance while refining tumor edges through multi-scale fusion [2].
@@
In complex models, dynamically adjusting the weights of individual loss functions during training is essential for effectively balancing them. Approaches like SoftAdapt dynamically adjust the weights of loss components, such as mean squared error (MSE) and Kullback-Leibler (KL) divergence, to optimize tasks like autoencoders and variational autoencoders. This ensures a more balanced learning process, preventing the model from overly focusing on any criterion. [3]. The dynamically weighted balanced (DWB) loss proposed by Ruwani et al. [4] improves accuracy in class-imbalanced datasets by adjusting weights based on class frequency and prediction difficulty. These dynamic weighting strategies are crucial in applications where class imbalance significantly impacts model performance.
@@
Another example of an adaptive loss function application is found in unsupervised image segmentation. Guermazi et al. [5] developed a dynamically weighted loss function that balances feature similarity and spatial continuity during training. The balance is automatically adapted based on cluster predictions, allowing more accurate and semantically meaningful segmentations. Similar dynamic approaches are applied in multiorgan segmentation, where Song et al. [6] introduced a dynamic loss weighting algorithm that adjusts the weight assigned to each organ based on its learning difficulty, ensuring more consistent performance across organs with varying sizes and complexities.
@@
The need for customized composite loss functions also extends to image fusion and classification tasks. Research such as FuseFormer [7] and S2F-Net [8] have introduced loss functions that balance pixel-level accuracy and structural similarity to ensure quality fusion results in infrared and visible image fusion. Finger vein recognition (FVR) also benefits from composite loss functions, where combining cosine softmax and triplet loss improves feature learning, generalization, and robustness [9]. In the context of ultrasound diagnosis systems, combining class activation mapping (CAM)-based loss and quantitative ultrasound (QUS)-based loss using fixed weights improves classification accuracy by enhancing lesion discrimination [10].
Meta-learning techniques have also been applied to dynamic loss weighting, as shown by Jiang et al. [11], who proposed a meta-learning loss function that adapts to label noise and class imbalance in biased datasets by dynamically adjusting loss weights based on learned margins. These approaches demonstrate the effectiveness of both composite and dynamic loss functions in enhancing performance across a wide range of deep learning applications.
@@
In this research, we introduce the dynamic memory fusion (DMF) framework, which builds upon these advancements by dynamically adjusting the weights of multiple base loss functions during training. Unlike fixed-weight approaches, the DMF framework leverages historical loss data to dynamically adjust the weighting of individual loss functions in real time. This adaptive approach ensures a balanced focus on different aspects of the learning task, preventing the model from becoming overly biased towards any single component. Our framework draws inspiration from previous studies on composite loss functions and dynamic weighting, offering a solution for multi-loss function optimization in deep learning.
@@ 2 Background
Adaptive loss functions have proven effective in segmentation tasks where visual detail is crucial. Chao et al. [18] introduced an adaptive composite loss function for a multi-field-of-view (FoV) visual saliency model in 360° image processing. This method dynamically adjusts the weights of Kullback-Leibler divergence (KLD), normalized scanpath saliency (NSS), and linear correlation coefficient (CC) based on their standard deviations during training. By doing so, the model optimizes performance by balancing key saliency features. Similarly, B et al. [19] developed a deep learning model for breast lesion segmentation in ultrasound images using a composite loss that combines binary cross-entropy (BCE) and boundary loss, demonstrating the efficacy of composite loss functions in medical image analysis.
@@ 3.1 Dynamic Memory Fusion Framework
$\mathcal{L}(y,\hat{y})=\sum_{i=1}^{N}w_{i}\mathcal{L}_{i}(y,\hat{y})+\gamma(t)\mathcal{L}_{a}(y,\hat{y})$ (1)
where $y$ and $\hat{y}$ refer to ground truth and predicted labels, respectively; $N$ represents the number of base loss functions, $\mathcal{L}_{i}$ are the individual loss functions, $w_{i}$ are adaptive weights dynamically adjust based on the historical data of each loss function, $\mathcal{L}_{a}$ denotes the auxiliary loss, which can be any task-specific function critical in early training stages, $\gamma$ is a scaling factor controlling the contribution of the auxiliary loss, and $t$ indicates the current training step.
@@ 4.2 Variance-Based Weighting
This normalization ensures that loss functions exhibiting higher variance are assigned greater weight, reflecting their increased contribution to the overall optimization process.
@@ 4.5 Comparison of Weighting Methods
The Bayesian-based weighting method integrates both prior knowledge and real-time data to update loss function weights dynamically. By assessing the likelihood of each loss function’s contribution to the overall performance, this method allows the model to combine predefined expectations, such as the relative importance of certain tasks, with observed performance data. This results in a more informed and adaptive training strategy, where the model balances accumulated learning with prior criteria, leading to a more refined and targeted optimization process.
@@ 6.2 Mean Intersection over Union (Mean IoU)
$\mathcal{L}_{IoU}=1-\frac{1}{C}\sum_{c=1}^{C}\frac{TP_{c}}{TP_{c}+FP_{c}+FN_{c}}$ (17)
where $TP_{c}$, $FP_{c}$, and $FN_{c}$ represent the true positives, false positives, and false negatives for class $c$, respectively, and $C$ is the number of classes.
@@ 6.3 Mean Dice Loss
$\mathcal{L}_{Dice}=1-\frac{1}{C}\sum_{c=1}^{C}\frac{2\times TP_{c}}{2\times TP_{c}+FP_{c}+FN_{c}}$ (18)
where $TP_{c}$, $FP_{c}$, and $FN_{c}$ correspond to the true positives, false positives, and false negatives for class $c$, and $C$ represents the number of classes.
@@ 6.4 Focal Loss
where $\alpha_{c}$ is a balancing factor for class $c$, $\gamma$, is a focusing parameter that adjusts the rate at which easy examples are down-weighted, $y_{ic}$ is the ground truth, and $\hat{y}_{ic}$ is the predicted probability.
@@ 6.5 Tversky Loss
where $\alpha$ and $\beta$ are hyperparameters that control the penalty for false positives and false negatives, respectively. This formulation allows the model to be more sensitive to either false positives or false negatives based on the specific task requirements.
7 Class Balanced Dice Loss
@@
class $\textit{class ratio}=\frac{P_{i}}{P}$ (21)
@@
where $P_{i}$ represents the number of pixels belonging to class $i$, and $P$ is the total number of pixels in the image. By subtracting the class ratio from 1, we ensure that classes with fewer pixels receive higher weights, balancing the impact of majority and minority classes on the final loss value. The core of the CB-Dice loss is computed similarly to the traditional dice score but includes class weights. The dice score for each class is computed using the true positives (TP), false positives (FP), and false negatives (FN), as follows:
@@ 10 Results
This section presents the results of applying the DMF framework with different auxiliary loss functions. Each experiment involved training models using three different weighting methods: variance-based, MAD-based, and Bayesian-based. To maintain fairness in comparisons, we utilized a basic U-Net model across all experiments to ensure that the methods, rather than model architectures, are being compared. While more advanced models could enhance performance, our goal is to isolate the benefits of the DMF framework itself. For every method, ten experiments were conducted with the same weight initialization and data distribution for each experiment, where 70% of the data was used for training, 15% for validation, and 15% for testing, ensuring consistent comparison. The evaluation metrics include dice, IoU, f1-score, precision, recall, and CB-Dice. The loss functions used in each experiment comprised categorical cross-entropy, mean IoU, and mean Dice as primary loss functions; additionally, Tversky, Focal, and CB-Dice were utilized as auxiliary loss functions. The demo source code for implementing the proposed methodology is openly available on GitHub at https://github.com/amingolnari/Demo-Dynamic-Memory-Fusion-Framework. To facilitate usage, readers can run the code directly in Google Colab.
@@
Overall, these results demonstrate that the choice of auxiliary loss function and the weighting method play critical roles in optimizing segmentation performance on each dataset.

### v2

@@ Abstract
[inst1]organization=Faculty of Electrical Engineering, Shahrood University of Technology, city=Shahrood, country=Iran
@@ 1 Introduction
Deep learning has revolutionized numerous domains, with medical imaging being one of the most transformative areas. Tasks such as segmentation, classification, and prediction have seen significant advancements due to sophisticated deep learning techniques. However, a key challenge remains, optimizing loss functions to guide model training effectively. Loss functions are critical in determining model performance, yet traditional approaches often rely on manually assigning fixed weights to multiple loss components; a process that is both labor-intensive and suboptimal in dynamic training environments. This issue is further compounded by class imbalance in medical imaging datasets, where underrepresented classes are frequently overlooked, leading to biased learning outcomes. Addressing these challenges requires an innovative approach that dynamically adjusts loss function weights to enhance model adaptability and performance.
This study addresses the pressing need for improved multi-loss function optimization in real-time scenarios, particularly in complex tasks like breast cancer segmentation. Current frameworks often struggle with adaptability when faced with evolving datasets. To overcome these limitations, we propose the Dynamic Memory Fusion (DMF) framework, which adaptively adjusts the weighting of multiple loss functions based on historical data. Additionally, we introduce a class-balanced dice loss function specifically designed to mitigate class imbalance, ensuring that underrepresented classes receive adequate attention during training. By tackling these challenges, our approach aims to provide a robust solution for real-time medical image segmentation, ultimately enhancing diagnostic accuracy and efficiency.
@@
Dynamic weighting of loss functions are crucial for optimizing performance across diverse tasks. This is especially important when addressing challenges such as class imbalance, boundary precision, and feature learning. For instance, GBE-Net [1] demonstrates the benefits of combining cross-entropy and dice loss to improve segmentation accuracy, particularly in preserving lesion boundaries. Similarly, MFMSNet [2] integrates dice and binary cross-entropy losses to address class imbalance while refining tumor edges through multi-scale fusion.
Despite these advances, many studies still rely on fixed-weight schemes for simplicity and consistency, such as those proposed in [3, 4, 5, 6, 7]. While these methods achieve reasonable performance by maintaining static weights throughout training, they often fail to adapt to varying complexities or imbalances within datasets. This limitation underscores the need for more flexible approaches.
@@
Dynamic weighting strategies have emerged as a promising solution. For example, SoftAdapt [8] dynamically adjusts the weights of loss components, such as mean squared error (MSE) and Kullback-Leibler Divergence (KLD), to optimize tasks like autoencoders and variational autoencoders. Similarly, the dynamically weighted balanced (DWB) loss proposed by Ruwani et al. [9] improves accuracy in class-imbalanced datasets by adjusting weights based on class frequency and prediction difficulty. These approaches highlight the importance of adaptability in achieving balanced learning.
@@
Dynamic weighting has also proven effective in specialized applications. In unsupervised image segmentation, Guermazi et al. [10] developed a dynamically weighted loss function that balances feature similarity and spatial continuity during training. Song et al. [11] introduced a similar approach for multiorgan segmentation, where the weight assigned to each organ is adjusted based on its learning difficulty, ensuring consistent performance across organs with varying sizes and complexities.
@@
Beyond segmentation, adaptive loss functions have been applied to tasks such as image fusion and classification. FuseFormer [12] and S2F-Net [13] balance pixel-level accuracy and structural similarity for infrared and visible image fusion. In finger vein recognition, combining cosine softmax and triplet loss improves feature learning and robustness [14]. For ultrasound diagnosis systems, fixed-weight combinations of CAM-based and QUS-based losses enhance lesion discrimination [15]. Meta-learning techniques, such as those proposed by Jiang et al. [16], further demonstrate the potential of dynamic loss weighting to adapt to label noise and class imbalance in biased datasets.
@@
Building on these advancements, the DMF framework introduces a novel approach to multi-loss function optimization. Unlike fixed-weight strategies, DMF leverages historical loss data to dynamically adjust the weighting of individual loss functions in real time. This ensures a balanced focus on different aspects of the learning task, preventing the model from becoming overly biased toward any single component. Our framework draws inspiration from prior work on composite loss functions and dynamic weighting but extends these concepts to provide a more adaptive and scalable solution for deep learning applications.
The contributions of this paper address fundamental challenges in medical image segmentation, including class imbalance and the manual tuning of multiple loss functions. Specifically, our key contributions are as follows:
Dynamic Memory Fusion Framework: We propose a novel framework for adaptive multi-loss function optimization, enabling real-time adjustment of loss weights based on historical data. Class-Balanced Loss Function: We introduce a new loss function designed to address class imbalance, prioritizing underrepresented classes and improving segmentation performance for minority classes. Comprehensive Evaluation: We conduct experiments on three breast ultrasound datasets (BUSI, BUSC, and BUS Synthetic), demonstrating the effectiveness of the DMF framework across various metrics. Open-Source Code: To ensure reproducibility and encourage further research, we provide an open-source implementation of the DMF framework. By addressing the limitations of existing approaches and introducing innovations in dynamic weighting and class balance, this work aims to advance the state of the art in medical image segmentation and beyond.
@@ 2 Background
Adaptive loss functions have proven effective in segmentation tasks where visual detail is crucial. Chao et al. [23] introduced an adaptive composite loss function for a multi-field-of-view (FoV) visual saliency model in 360° image processing. This method dynamically adjusts the weights of KLD, normalized scanpath saliency (NSS), and linear correlation coefficient (CC) based on their standard deviations during training. By doing so, the model optimizes performance by balancing key saliency features. Similarly, B et al. [24] developed a deep learning model for breast lesion segmentation in ultrasound images using a composite loss that combines binary cross-entropy (BCE) and boundary loss, demonstrating the efficacy of composite loss functions in medical image analysis.
@@ 3.1 Dynamic Memory Fusion Framework
$\mathcal{L}(y,\hat{y},t)=\sum_{i=1}^{N}w_{i}\mathcal{L}_{i}(y,\hat{y})+\gamma(t)\mathcal{L}_{a}(y,\hat{y})$ (1)
where $y$ and $\hat{y}$ refer to ground truth and predicted labels, respectively; $t$ refers to the cycle completed (which could be a batch or epoch), $N$ represents the number of base loss functions, $\mathcal{L}_{i}$ are the individual loss functions, $w_{i}$ are adaptive weights dynamically adjust based on the historical data of each loss function, $\mathcal{L}_{a}$ denotes the auxiliary loss, which can be any task-specific function critical in early training stages, $\gamma$ is a scaling factor controlling the contribution of the auxiliary loss, and $t$ indicates the current training step.
@@ 4.2 Variance-Based Weighting
This normalization ensures that loss functions exhibiting higher variance are assigned proportionally greater weight within the interval [0,1][0,1][ 0 , 1 ], reflecting their increased contribution to the overall optimization process.
@@ 4.5 Comparison of Weighting Methods
The Bayesian-based weighting method integrates both prior knowledge and real-time data to update loss function weights dynamically. By assessing the likelihood of each loss function’s contribution to the overall performance, this method allows the model to combine predefined expectations, such as the relative importance of certain tasks, with observed performance data. This results in a more informed and adaptive training strategy, where the model balances accumulated learning with prior criteria, leading to a more refined and targeted optimization process. Summarized comparative analysis is as follow:
  • 1.
    Variance-Based Weighting: Focuses on loss functions with higher variability, assuming they represent challenging aspects of the task.
  • 2.
    MAD-Based Weighting: Prioritizes stability by emphasizing loss functions with smaller deviations from their median values, making it robust to outliers.
  • 3.
    Bayesian-Based Weighting: Integrates prior knowledge with observed data, enabling informed and adaptive weight adjustments.
@@ 6.2 Mean Intersection over Union (Mean IoU)
$\mathcal{L}_{IoU}=1-\frac{1}{N_{c}}\sum_{c=1}^{N_{c}}\frac{TP_{c}}{TP_{c}+FP_{c}+FN_{c}}$ (17)
where $c$ represents the class and $N_{c}$ is the number of classes.
@@ 6.3 Mean Dice Loss
$\mathcal{L}_{Dice}=1-\frac{1}{N_{c}}\sum_{c=1}^{N_{c}}\frac{2\times TP_{c}}{2\times TP_{c}+FP_{c}+FN_{c}}$ (18)
@@ 6.4 Focal Loss
where $\alpha_{c}$ is a balancing factor for class $c$, $\gamma$, is a focusing parameter that adjusts the rate at which easy examples are down-weighted, $y_{ic}$ is the ground truth, and $\hat{y}_{ic}$ is the predicted probability. The values for the balancing factor ($\alpha_{c}$ = 0.25) and focusing parameter ($\gamma$ = 2) in Focal Loss were empirically optimized to address class imbalance and down-weight easy examples, respectively, based on prior literature [34] and experimental validation.
@@ 6.5 Tversky Loss
where $\alpha$ and $\beta$ are hyperparameters that control the penalty for FPs and FNs, respectively. This formulation allows the model to be more sensitive to FP or FN based on the specific task requirements. The Tversky Loss parameters were empirically set to 0.7 for $\alpha$ and 0.3 for $\beta$ to balance their trade-off according to task requirements.
6.6 Class Balanced Dice Loss
@@
$\textit{class ratio}=\frac{P_{i}}{P}$ (21)
@@
where $P_{i}$ represents the number of pixels belonging to class $i$, and $P$ is the total number of pixels in the image. By subtracting the class ratio from 1 (Equation 25), we ensure that classes with fewer pixels receive higher weights, balancing the impact of majority and minority classes on the final loss value. The core of the CB-Dice loss is computed similarly to the traditional dice score but includes class weights. The dice score for each class is computed using the TP, FP, and FN, as follows:
@@
The initial weights for base loss functions were set equally to $1/N$, where $N$ represents the number of base loss functions, while the scaling factor ($\gamma_{0}$) for the auxiliary loss was experimentally determined to be 10, with a decay rate ($\tau$) of 0.05 applied after each batch, ensuring its gradual reduction during training. Figure 1 shows how the auxiliary loss scaling factor decreases gradually throughout the training process.
Figure 1: Weight decay of the auxiliary loss using scaling factor ($\gamma_{t}$) during training.
Table 1: Comparison of Loss Functions
Loss Function Technical Details Best Use Cases Strengths Limitations
Categorical Cross-Entropy Measures label-prediction mismatch Balanced class tasks Simple, widely used Sensitive to imbalance
Mean IoU Evaluates overlap between masks Segmentation with overlap focus Focuses on spatial accuracy Unstable for small classes
Mean Dice Loss Derived from Dice coefficient Boundary precision tasks Handles imbalance better Sensitive to noise
Focal Loss Modulates cross-entropy for hard examples Imbalanced datasets Reduces easy example impact Needs hyperparameter tuning
Tversky Loss Balances false positives/negatives Tasks needing FP/FN control Flexible trade-offs Hyperparameter sensitive
CB-Dice Loss Adds class weights to Dice loss Severe imbalance tasks Prioritizes minority classes Computationally complex
@@ 9 Results
This section presents the results of applying the DMF framework with different auxiliary loss functions. Each experiment involved training models using three different weighting methods: variance-based, MAD-based, and Bayesian-based. To maintain fairness in comparisons, we utilized a basic U-Net model across all experiments to ensure that the methods, rather than model architectures, are being compared. While more advanced models could enhance performance, our goal is to isolate the benefits of the DMF framework itself. As shown in Figure 2, The U-Net architecture used in this study follows a standard encoder-decoder structure, with convolutional blocks (ConvBlock) for feature extraction and downsampling, followed by transposed convolutional blocks (ConvTransposeBlock) for upsampling and reconstruction. Each block incorporates batch normalization, ReLU activation, and dropout to enhance training stability and generalization, while skip connections between corresponding encoder and decoder layers ensure precise localization of features.
Figure 2: Architecture of the U-Net model used in this study.
For every method, ten experiments were conducted with the same weight initialization and data distribution for each experiment, where 70% of the data was used for training, 15% for validation, and 15% for testing, ensuring consistent comparison. In our experiments, we used the Adam optimizer with a learning rate of $1\times 10^{-3}$ and a batch size of 64, implementing an algorithm to save the optimal weights based on validation dataset, which was originally proposed in our previous work [39], to prevent overfitting, and trained all models for the same number of epochs without showing signs of overfitting. To address overfitting concerns, in each experiment we applied random shuffling before splitting the dataset into training, validation, and test subsets. This ensured diverse and balanced data exposure across iterations. Table 2 provides a detailed summary of the experimental configuration.
Table 2: Experimental Settings Summary – Random Shuffling Method
Parameter Value / Description
Experiments per dataset 10 times
Weight initialization Identical initial weights across each experiment
Data splitting strategy Random shuffling followed by stratified split
Train/Val/Test ratios 70%/15%/15%
Augmentation Enabled with configurable probabilities (real-time)
Batch size 64
Learning rate $1\times 10^{-3}$
Epochs per run Fixed number of epochs (set to 100)
Evaluation metrics Dice, IoU, F1-Score, Precision, Recall, CB-Dice
Additionally, regularization techniques (batch normalization, 25% dropout, and L1L2 regularization), and data augmentation methods such as random flipping, rotations, and shifts used for enhancing model generalization and reducing the risk of overfitting during training. Instead of early stopping, as mentioned above, an algorithm tracked optimal weights based on validation dataset, while learning curves confirmed no divergence between training and validation losses. The evaluation metrics include dice, IoU, f1-score, precision, recall, and CB-Dice. The loss functions used in each experiment comprised categorical cross-entropy, mean IoU, and mean Dice as primary loss functions; additionally, Tversky, Focal, and CB-Dice were utilized as auxiliary loss functions.
To improve training stability, particularly in the presence of class imbalance, we combined the Mean Dice Loss and Mean IoU Loss as complementary objectives. While these metrics are mathematically related, their gradient behavior during optimization differs, offering a more balanced learning signal that reduces oscillations and improves convergence. This dual-loss strategy enhances robustness by simultaneously encouraging high overlap between predicted and ground truth regions (via Dice) and penalizing spatial mismatches more strictly (via IoU).
The demo source code for implementing the proposed methodology is openly available on GitHub at https://github.com/amingolnari/Demo-Dynamic-Memory-Fusion-Framework. To facilitate usage, readers can run the code directly in Google Colab.
@@
In summary, Variance-based weighting with CB-Dice loss consistently outperforms other configurations across all datasets, while MAD-based weighting shows strong precision but slightly lower recall on the BUSC dataset, and Bayesian-based weighting performs well on synthetic data but struggles with real-world variability.
@@
Overall, these results demonstrate that the choice of auxiliary loss function and the weighting method play critical roles in optimizing segmentation performance on each dataset. As illustrated in Figure 6, the training and validation loss curves ($\mathcal{L}(y,\hat{y},t)$) demonstrate a consistent decline without any significant divergence, indicating that the model does not suffer from overfitting during training.
Figure 6: Training and validation loss curves across epochs.
@@
To further validate the generalization capability of our method and ensure its reliability across different data distributions, we conducted an additional evaluation using 5-fold cross-validation. In our previous experimental setup, we trained the U-Net 10 times on each dataset with identical initial weights but different train/validation/test splits, achieved through random shuffling. This strategy ensured diverse exposure to the data and provided a stable estimate of model performance. As an extension, we have also applied k-fold cross-validation to further assess the consistency of our results. The fold-averaged performance metrics showed close alignment with those obtained from the repeated random splits, confirming that our findings are not biased by any specific data partitioning strategy and can be reliably generalized. Table 6 provides a detailed summary of the experimental configuration for the k-fold cross-validation procedure, whereas Table 7 outlines the main results obtained from the evaluation.
Table 6: Experimental Settings Summary – k-Fold Cross-Validation
Parameter Value / Description
Number of folds 5 folds
Dataset coverage Each sample used exactly once per fold as test set
Weight initialization Independent initialization per fold
Data splitting strategy Stratified k-fold (train/val/test)
Train/Val/Test ratios 70%/15%/15%
Augmentation Enabled with configurable probabilities (real-time)
Batch size 64
Learning rate $1\times 10^{-3}$
Epochs per run Fixed number of epochs (set to 100)
Evaluation metrics Dice, IoU, F1-Score, Precision, Recall, CB-Dice
Table 7: Performance metrics obtained through 5-fold cross-validation across all datasets using the CB-Dice loss as the auxiliary loss function. Reported values represent the mean performance across the five folds.
Dataset Method Dice IoU F1-score Precision Recall CB-Dice
BUSI Variance 94.98 91.57 95.25 95.89 94.62 92.83
MAD 95.16 90.83 95.02 95.29 94.75 92.12
Bayesian 92.73 87.48 93.28 91.37 95.27 88.87
BUSC Variance 95.53 91.49 95.75 94.69 96.83 93.47
MAD 94.32 89.64 94.48 94.30 94.67 92.65
Bayesian 95.16 89.85 95.07 94.26 95.89 93.37
BUS Variance 95.84 91.69 95.47 95.48 95.46 93.87
MAD 94.89 91.37 94.35 92.41 96.38 93.47
Bayesian 92.98 87.26 93.60 90.65 96.73 89.76
Table 8 provides a detailed comparison of several existing approaches to dynamic and adaptive loss function weighting, highlighting their respective advantages, limitations, and associated performance outcomes.
Table 8: Summary of Dynamic Weighting Methods for Loss Functions
Research Short Description Pros Cons Results
[23] Dynamically balances KLD, CC, and NSS metrics during training based on standard deviations Adapts to dataset characteristics; improves performance across metrics Requires tuning of initial weights Outperforms state-of-the-art methods on multiple datasets
[26] Adjusts cross-entropy weights dynamically using batch and global data frequencies Handles class imbalance effectively; improves segmentation accuracy Hard truncation may discard useful information Achieves better performance on imbalanced datasets
[10] Dynamically adjusts weights between feature similarity and spatial continuity Balances focus between similarity and continuity; adapts to training dynamics Requires careful tuning of parameters Improves semantic segmentation details
[33] Dynamically adjusts shape parameter $\alpha$ to transition between traditional loss functions Eliminates manual tuning; adapts to varying data May introduce instability if $\alpha$ changes too rapidly Performs well across diverse tasks and datasets
[31] Dynamically adjusts weights for energy, forces, and stress during training Reduces need for manual tuning; balances contributions Sensitive to initial conditions and learning rates Improves robustness in physics-informed neural networks
[16] Dynamically reduces impact of noisy or hard-to-learn samples Improves robustness to noise; enhances generalization May underemphasize difficult samples Achieves state-of-the-art performance on noisy datasets
[11] Uses an autoencoder to assign larger weights to difficult organs Promotes consistent performance across organs; eliminates manual tuning Adds computational overhead Outperforms baselines and state-of-the-art methods on multiorgan segmentation
[28] Dynamically adjusts weights based on error rates observed after each training round Effectively handles class imbalance; improves prediction accuracy Requires additional computation for error rate calculations Improves sentiment classification accuracy
[29] Combines L-softmax with adaptive weighting based on inverse class frequencies Addresses class imbalance; reduces overfitting Requires careful tuning of margin parameters Enhances performance on imbalanced datasets
[30] Dynamically adjusts weights based on coefficient of variation of each loss component Eliminates manual tuning; adapts to training dynamics Struggles with highly volatile loss components Improves balanced performance across multiple tasks
[27] Dynamically adjusts class weights using Ordered Weighted Averaging (OWA) operators Versatile and easy to implement; addresses class imbalance Requires tuning of aggregation parameters Outperforms focal loss and cross-entropy on classification tasks
[32] Dynamically adjusts weights of physics and boundary losses Improves robustness in PINNs; reduces sensitivity to weight choices Requires careful initialization of adaptive parameters Enhances performance in physics-informed neural networks
[8] Dynamically updates weights of multi-part loss functions based on live performance metrics Fast and easy to implement; compatible with any gradient-based optimizer Sensitive to hyperparameter $\beta$ Improves performance in multi-objective optimization tasks
Ours Dynamically adjusts weights using Variance, MAD, and Bayesian-based methods Eliminates manual tuning; Adapts to dynamic scenarios; Flexible strategies Requires careful selection of weighting method Achieves superior performance on breast ultrasound image segmentation